solve the inequality
4k^3-6k^2(< or =)4k

Answer by RAY100(1637)

4k^3 -6k^2 < = 4k
.
lets solve as equality first
.
4k^3 -6k^2 = 4k
.
4k^3 -6k^2 -4k =0
.
2k ( 2k^2 -3k -2) = 0
.
2k (2k+1)(k-2) =0
.
2k =0,,,,k=0
.
2k+1 =0,,,,k= (-1/2)
.
k-2 = 0,,,,k=2
.
Now on number line we have,,, (-1/2),,,,(0),,,,,(2)
.
Now lets check if numbers before and after each of these satisfy original inequality
.
for x< (-1/2),,,,let x = -1,,,,-10 <= -4,,,,ok
.
for (-1/2) < x < 0,,,let x= (-.25),,,-.3125 <=-1,,,,not true
.
for 0 less than x less than 2,,,,,let x=1,,,,-2 <= 4,,,,ok
.
for x>2,,,,let x= 3,,,,,54<=12,,,,,not true
.
Therefore,,,,,x<=(-1/2),,,,,0<=x<=2,,,,,answer